# Co-Training Vision Language Models for Remote Sensing Multi-task Learning

### v1

@@ I Introduction
EARTH observation systems have acquired extensive remote sensing (RS) data, necessitating the development of automated RS image interpretation techniques [GRSM22review]. The emergence of artificial general intelligence has inspired researchers in the RS community to develop versatile agents capable of performing multiple tasks, such as scene classification, visual question answering, and object detection [zhou2024vlgfm].
@@
Most RS image processing methods typically train a specifically-designed model on isolated datasets to achieve optimal performance on individual tasks. Due to the heterogeneity of data and model architecture, developing a unified model capable of handling multiple RS tasks, i.e., multi-task learning (MTL), remains challenging [mtl_survey_2022].
@@
MTL provides several advantages for RS applications. First, a single MTL model with shared parameters can handle multiple tasks at once, unlike traditional task-specific models, which is closer to human perception. Second, by sharing knowledge across tasks, MTL mitigates the shortage of annotated data and reduces overfitting on individual tasks. Third, MTL learns joint representations that capture correlations among tasks, improving generalization. RS foundation models also benefit by obtaining consistent representations through pre-training on upstream tasks and fine-tuning on various downstream tasks. Overall, MTL helps advance RS foundation models by expanding pre-training tasks and enhancing cross-task learning [li2024co].
@@
Transformer [Transformer] has demonstrated remarkable flexibility and generalization capabilities across various domains, including computer vision [ViT], natural language processing, speech processing, and remote sensing data analysis [TRD]. This progress has brought the goal of a unified multimodal and multi-task architecture increasingly within reach [han2023onellm]. Consequently, vision language models (VLMs), that bridge the gap between the two modalities by learning from vast amounts of paired [clip] and interleaved image-text data [li2024omnicorpus], have been proposed and become the most commonly adopted foundation model paradigm in the multimodal domain [flamingo].
@@
In this study, we focus on generative VLMs, also named multimodal language models. These models are typically constructed upon vision and language foundation models, enabling them to process visual inputs and effectively interpret textual instructions. By harnessing the capabilities of powerful pre-trained foundation models and leveraging a versatile text interface, VLMs are positioned as a crucial element in the progression toward the unified MTL [internvl].
@@
We consider that VLMs represent an ideal paradigm for RS MTL. Firstly, the textual interface of VLMs provides a unified representation for diverse task objectives, because the outputs of different RS tasks, such as classification, grounding, captioning, or question answering, can all be expressed in text form. Secondly, instruction tuning has demonstrated that VLMs can generalize beyond the tasks seen during training [llava], enabling them to handle novel or composite tasks through in-context learning [flamingo]. Finally, with sufficiently strong foundational capabilities, VLMs offer the potential to evolve toward more autonomous RS agents, where task reasoning and workflow design can be accomplished within a single, coherent framework.
@@
Early RS MTL approaches were typically designed for multiple pure-vision tasks [li2024co, MTP, li2024sm3det], such as classification, segmentation, and detection. These methods generally adopt a shared feature extraction backbone with task-specific output heads. With carefully crafted training strategies, their performance on individual benchmarks is comparable to that of expert models trained on the specific dataset. However, they suffer from limited scalability and architectural rigidity. As the number of tasks increases, the heterogeneous design of multiple heads makes optimization increasingly difficult and less robust. Consequently, this paradigm struggles to scale up, resulting in insufficient model generalization. Nevertheless, when deployed on resource-constrained platforms such as satellites, this kind of MTL model remains highly valuable for its computational and storage efficiency.
As general-purpose VLMs increasingly exhibit early signs of the universal model, they have emerged as a scalable paradigm for MTL. In the RS domain, several studies have explored VLM-based MTL. However, their investigations into unified and generalizable paradigms remain limited: The regular VLMs focus primarily on language-centric description tasks such as image captioning and visual question answering, where text descriptions are synthesized for RS images to enable semantic understanding [vhm, geochat, lhrsbot]. Others extend VLMs to purely visual tasks such as visual grounding and object detection, leveraging the flexible language interface of VLMs to learn from abundant localization annotations and achieve precise detection capabilities [geoground, lmmrotate]. In addition, several approaches target ultra-high-resolution (UHR) RS image reasoning, often employing token pruning to alleviate the computational burden caused by extremely large inputs [lrsvqa, geollava8k].
These studies collectively highlight the great potential of VLMs for RS MTL, yet each remains constrained within a limited scope. As shown in Fig. 1, the first four types of works focus mainly on tasks involving regular images (images with regular resolutions) [vhm, geochat, lhrsbot, geoground, lmmrotate], whereas the fifth is specialized for UHR scenarios [lrsvqa, geollava8k]. The detection VLMs [lmmrotate] and grounded VLMs [geoground] excel at spatial grounding but pays little attention to semantic understanding, while the regular VLMs [vhm, geochat, lhrsbot] rarely explore the crucial object detection capabilities which is essential for RS image analysis. Hence, a unified framework that addresses these limitations in an integrated MTL setting is still lacking.
@@
Finally, previous VLMs have shown limited capability in object detection tasks. They either perform only sparse visual grounding [geochat, geoground], provide detection results for a single category [earthgpt], or are evaluated leniently under low IoU thresholds [luo2024skysensegpt]. However, aerial detection is particularly challenging due to issues such as dense object distribution, which places high demands on the visual input resolution and output sequence length of MLLMs. Moreover, VLMs cannot directly output the confidence scores of predicted objects, making it difficult to fairly compare them with traditional models using commonly-used evaluation metrics.
@@
This manuscript is an extended and improved version of our conference paper [lmmrotate] published in IGARSS 2025, which only investigate VLMs for detection tasks. The autoregressive object detection scheme in Section III-E is primarily derived from the conference version. Building upon it, we not only refine detection details in Section III-E1 but also further upgrade the VLMs with unified multi-task learning, accompanied by additional methods, models, and experimental results.
@@ II-A General-purpose Vision Language Models
Early works such as VisualGPT [visualgpt], BLIP-2 [blip2], and Flamingo [flamingo] explored different ways of integrating visual features with large language models or training joint image-text encoders, showing improved multimodal reasoning and understanding.
Later instruction-tuned frameworks, including LLaVA [llava], MiniGPT-4 [zhu2024minigpt], and InstructBLIP [InstructBlip], further enhanced interactive comprehension by fine-tuning LLMs with visual–text instructions. Lightweight adaptation methods such as LLaMA-Adapter V2 [lamaadapterv2] and SPHINX [sphinx] improved efficiency through visual adapters and zero-shot attention fusion, reducing the cost of multimodal alignment.
In parallel, the VisionLLM series [VisionLLM, visionllm-v2] unified vision-centric tasks under the LLM paradigm, enabling open-ended reasoning over diverse visual inputs. More recent large-scale MLLMs, including PaLI-X [palix], MiMo-VL [mimovl], InternVL [internvl], CogVLM [cogvlm], and the Qwen-VL series [qwen2vl, qwen25vl], further integrate vision and language through end-to-end pretraining on massive multimodal data and scalable architectures. These models show improved visual grounding, OCR, and cross-domain reasoning, representing a shift from adapter-based fusion toward deeply coupled vision-language modeling. Collectively, these advances lay the foundation for adapting MLLMs to specialized domains such as remote sensing, where complex visual semantics and open-ended reasoning are required.
@@ II-B Remote Sensing Vision Language Models
Recently, integrating vision–language models into remote sensing (RS) has attracted growing attention, giving rise to several domain-specific VLMs. GeoChat [geochat] pioneered this direction as the first RS-oriented VLM, addressing multiple optical imagery tasks via conversational interaction. EarthGPT [earthgpt] introduced a unified multimodal framework for multi-source RS data and diverse vision-language tasks. LHRS-Bot [lhrsbot] leverages multi-level vision-language alignment and curriculum learning for RS image understanding.
Beyond static image interpretation, recent works explore temporal and fine-grained understanding. TEOChat [irvin2024teochat] supports temporal Earth observation imagery and instruction following over sequential frames. SkySenseGPT [luo2024skysensegpt] extends instruction tuning to fine-grained RS comprehension, achieving strong performance on public datasets and complex comprehension tasks. VHM [vhm] demonstrates capabilities on tasks such as building vectorization, multi-label classification, and honest question answering.
@@ II-C Remote Sensing Multi-task Learning
One line of research focuses on leveraging shared representations for synergistic task pairing. For example, several studies [MTSCD-Net, stgnet, SMNet, mcenet] combine semantic segmentation with change detection from bi-temporal imagery, showing that joint learning enhances feature sharing and reduces redundancy. Another direction jointly models geometric and semantic information, such as height estimation with semantic segmentation [lu2022multi, carvalho2019multitask], demonstrating gains over single-task baselines.
Beyond specific task pairs, generalized MTL frameworks have been proposed. RSCoTr [li2024co] performs classification, segmentation, and detection simultaneously, illustrating the potential of unified RS analysis. Large-scale datasets like SatlasPretrain [Bastani_2023_ICCV] with multiple annotation modalities facilitate advanced MTL model development. SM3Det [li2024sm3det] uses a mixture-of-experts structure for multi-modal detection of horizontal and rotated bounding boxes. EarthDial [soni2025earthdial] leverages multiple multi-task decoders to transfer knowledge across diverse tasks, enriching shared feature learning.
@@ III-B1 Data Acquisition
In this work, the dataset was curated through three sequential stages. Initially, we collected data by following the data recipes of several representative open-source vision–language models. Specifically, we adopted the description–related subsets from the instruction tuning data of VHM [vhm] and GeoChat [geochat], which cover tasks such as image classification, captioning, and visual question answering. These tasks can be further decomposed into subtasks, including modality recognition and resolution estimation. The refGeo [geoground] dataset was employed as the main grounding data source, while temporal multi-image data were drawn from TEOChatlas [irvin2024teochat]. To prevent degradation of the model’s general reasoning ability during continued training, we also incorporated a subset of general-purpose data sampled from LLaVA-OneVision’s recipe [li2025llavaonevision], including chart interpretation, optical character recognition, and so on. By following these open-source data recipes, we indirectly surveyed and integrated diverse data sources.
Subsequently, we analyzed the limitations of the collected data and expanded the dataset using task-specific training set. We observed that existing RS VLMs rarely address object detection, which is crucial for fine-grained perception in RS. To fill this gap, we incorporated the DOTA-v1.0 dataset [dota], thereby enriching the model’s detection-related learning capabilities.
@@ III-B2 Data Processing and Integrating
We first removed all task descriptors, such as the “[grounding]”, “[refer]”, and “[identify]” tags used in previous works [geochat, vhm]. These descriptors tag the specific tasks. However, in open-world scenarios or novel tasks, instructions are typically expressed in natural language rather than through fixed descriptor tokens. Therefore, we replaced these descriptors with natural language prompts to better align with the real-world usage.
@@
A unified data format was further established to standardize the integration. Conversational data followed the messages structure defined by OpenAI, object detection data were formatted according to the COCO convention, and grounding data adhered to the refGeo [geoground] schema.
@@ III-B3 Data Loading and Weighting
We argue that the model should not rely solely on predefined prompts from the training stage. To enhance robustness, multiple agent prompts were designed for certain tasks, and one was randomly selected during training. For grounding and detection data, a unified formatting scheme was adopted. We also incorporated the JSON-based output format used in Qwen2.5-VL [qwen25vl], accompanied by specific prompts, and randomly switched between standard and JSON outputs during training. In addition, a synonym replacement module was implemented to randomly substitute words with their synonyms, improving the model’s linguistic generalization. Standard data augmentation techniques, such as random resizing, were also applied to enhance multi-scale learning.
@@ III-C Dynamic Resolution Strategy
Most existing RS VLMs (such as GeoChat [geochat], VHM [vhm], and GeoGround [geoground]) have the only fixed square input shape (such as $336\times 336$ or $504\times 504$). For each input image, they first pad the image to a square with zeros on the right or bottom, and then resize it to the input shape. Additionally, LRS-VQA [lrsvqa] and GeoLLaVA-8k [geollava8k] scale the input size to $2k\times 2k$ and $8k\times 8k$, respectively. They first cut a UHR RS image into slices of the fixed size and encode them into visual tokens. Then they prune the tokens to an amount comparable to the normal cases. In total, they pre-process the images only on a fixed shape or a small set of image shapes.
@@ III-C1 Full-scale Visual Input
The native resolution scheme in Qwen2-VL [qwen2vl] inspire us to advance RS VLMs to accept inputs of arbitrary shapes. As shown in Fig. 3, let $H$ and $W$ indicate the height and width of a given RS image. $L_{patch}$ is the patch length corresponding to each visual token from the vision encoder. They first calculate the tightest shape that can wrap the input image by
@@ III-C2 Scalable Bounding Boxes
However, existing RS VLMs often suffer from a mismatch between the coordinate resolution and the input image resolution. For instance, GeoChat [geochat] processes images at a fixed resolution of $\times504$, but its coordinate resolution is only $\times100$, leading to a fivefold loss in localization precision and poor performance on small objects. Conversely, GeoGround [geoground] employs a $\times336$ input resolution but defines coordinates at a much higher $\times1000$ scale, resulting in more than half of the coordinate space being unused and excessive localization precision.

### v2

@@ I Introduction
EARTH observation systems have acquired extensive remote sensing (RS) data, necessitating the development of automated RS image interpretation techniques [72]. The emergence of artificial general intelligence has inspired researchers in the RS community to develop versatile agents capable of performing multiple tasks, such as scene classification, visual question answering, and object detection [77].
@@
Most RS image processing methods typically train a specifically-designed model on isolated datasets to achieve optimal performance on individual tasks. Due to the heterogeneity of data and model architecture, developing a unified model capable of handling multiple RS tasks, i.e., multi-task learning (MTL), remains challenging [76].
@@
MTL provides several advantages for RS applications. First, a single MTL model with shared parameters can handle multiple tasks at once, unlike traditional task-specific models, which is closer to human perception. Second, by sharing knowledge across tasks, MTL mitigates the shortage of annotated data and reduces overfitting on individual tasks. Third, MTL learns joint representations that capture correlations among tasks, improving generalization. RS foundation models also benefit by obtaining consistent representations through pre-training on upstream tasks and fine-tuning on various downstream tasks. Overall, MTL helps advance RS foundation models by expanding pre-training tasks and enhancing cross-task learning [26].
@@
Transformer [53] has demonstrated remarkable flexibility and generalization capabilities across various domains, including computer vision [15], natural language processing, speech processing, and remote sensing data analysis [28]. This progress has brought the goal of a unified multimodal and multi-task architecture increasingly within reach [18]. Consequently, vision language models (VLMs), that bridge the gap between the two modalities by learning from vast amounts of paired [47] and interleaved image-text data [29], have been proposed and become the most commonly adopted foundation model paradigm in the multimodal domain [1].
@@
In this study, we focus on generative VLMs, also named multimodal language models. These models are typically constructed upon vision and language foundation models, enabling them to process visual inputs and effectively interpret textual instructions. By harnessing the capabilities of powerful pre-trained foundation models and leveraging a versatile text interface, VLMs are positioned as a crucial element in the progression toward the unified MTL [9].
@@
We consider that VLMs represent an ideal paradigm for RS MTL. Firstly, the textual interface of VLMs provides a unified representation for diverse task objectives, because the outputs of different RS tasks, such as classification, grounding, captioning, or question answering, can all be expressed in text form. Secondly, instruction tuning has demonstrated that VLMs can generalize beyond the tasks seen during training [38], enabling them to handle novel or composite tasks through in-context learning [1]. Finally, with sufficiently strong foundational capabilities, VLMs offer the potential to evolve toward more autonomous RS agents, where task reasoning and workflow design can be accomplished within a single, coherent framework.
@@
Early RS MTL approaches were typically designed for multiple pure-vision tasks [26, 54, 33], such as classification, segmentation, and detection. These methods generally adopt a shared feature extraction backbone with task-specific output heads. With carefully crafted training strategies, their performance on individual benchmarks is comparable to that of expert models trained on the specific dataset. However, they suffer from limited scalability and architectural rigidity. As the number of tasks increases, the heterogeneous design of multiple heads makes optimization increasingly difficult and less robust. Consequently, this paradigm struggles to scale up, resulting in insufficient model generalization. Nevertheless, when deployed on resource-constrained platforms such as satellites, this kind of MTL model remains highly valuable for its computational and storage efficiency.
As general-purpose VLMs increasingly exhibit early signs of the universal model, they have emerged as a scalable paradigm for MTL. In the RS domain, several studies have explored VLM-based MTL. However, their investigations into unified and generalizable paradigms remain limited: The regular VLMs focus primarily on language-centric description tasks such as image captioning and visual question answering, where text descriptions are synthesized for RS images to enable semantic understanding [46, 22, 43]. Others extend VLMs to purely visual tasks such as visual grounding and object detection, leveraging the flexible language interface of VLMs to learn from abundant localization annotations and achieve precise detection capabilities [78, 27]. In addition, several approaches target ultra-high-resolution (UHR) RS image reasoning, often employing token pruning to alleviate the computational burden caused by extremely large inputs [42, 55].
These studies collectively highlight the great potential of VLMs for RS MTL, yet each remains constrained within a limited scope. As shown in Fig. 1, the first four types of works focus mainly on tasks involving regular images (images with regular resolutions) [46, 22, 43, 78, 27], whereas the fifth is specialized for UHR scenarios [42, 55]. The detection VLMs [27] and grounded VLMs [78] excel at spatial grounding but pays little attention to semantic understanding, while the regular VLMs [46, 22, 43] rarely explore the crucial object detection capabilities which is essential for RS image analysis. Hence, a unified framework that addresses these limitations in an integrated MTL setting is still lacking.
@@
Finally, previous VLMs have shown limited capability in object detection tasks. They either perform only sparse visual grounding [22, 78], provide detection results for a single category [74], or are evaluated leniently under low IoU thresholds [41]. However, aerial detection is particularly challenging due to issues such as dense object distribution, which places high demands on the visual input resolution and output sequence length of MLLMs. Moreover, VLMs cannot directly output the confidence scores of predicted objects, making it difficult to fairly compare them with traditional models using commonly-used evaluation metrics.
@@
This manuscript is an extended and improved version of our conference paper [27] published in IGARSS 2025, which only investigate VLMs for detection tasks. The autoregressive object detection scheme in Section III-E is primarily derived from the conference version. Building upon it, we not only refine detection details in Section III-E1 but also further upgrade the VLMs with unified multi-task learning, accompanied by additional methods, models, and experimental results.
@@ II-A General-purpose Vision Language Models
Early works such as VisualGPT [5], BLIP-2 [25], and Flamingo [1] explored different ways of integrating visual features with large language models or training joint image-text encoders, showing improved multimodal reasoning and understanding.
Later instruction-tuned frameworks, including LLaVA [38], MiniGPT-4 [81], and InstructBLIP [13], further enhanced interactive comprehension by fine-tuning LLMs with visual–text instructions. Lightweight adaptation methods such as LLaMA-Adapter V2 [16] and SPHINX [36] improved efficiency through visual adapters and zero-shot attention fusion, reducing the cost of multimodal alignment.
In parallel, the VisionLLM series [60, 62] unified vision-centric tasks under the LLM paradigm, enabling open-ended reasoning over diverse visual inputs. More recent large-scale MLLMs, including PaLI-X [7], MiMo-VL [70], InternVL [9], CogVLM [57], and the Qwen-VL series [56, 2], further integrate vision and language through end-to-end pretraining on massive multimodal data and scalable architectures. These models show improved visual grounding, OCR, and cross-domain reasoning, representing a shift from adapter-based fusion toward deeply coupled vision-language modeling. Collectively, these advances lay the foundation for adapting MLLMs to specialized domains such as remote sensing, where complex visual semantics and open-ended reasoning are required.
@@ II-B Remote Sensing Vision Language Models
Recently, integrating vision–language models into remote sensing (RS) has attracted growing attention, giving rise to several domain-specific VLMs. GeoChat [22] pioneered this direction as the first RS-oriented VLM, addressing multiple optical imagery tasks via conversational interaction. EarthGPT [74] introduced a unified multimodal framework for multi-source RS data and diverse vision-language tasks. LHRS-Bot [43] leverages multi-level vision-language alignment and curriculum learning for RS image understanding.
Beyond static image interpretation, recent works explore temporal and fine-grained understanding. TEOChat [20] supports temporal Earth observation imagery and instruction following over sequential frames. SkySenseGPT [41] extends instruction tuning to fine-grained RS comprehension, achieving strong performance on public datasets and complex comprehension tasks. VHM [46] demonstrates capabilities on tasks such as building vectorization, multi-label classification, and honest question answering.
@@ II-C Remote Sensing Multi-task Learning
One line of research focuses on leveraging shared representations for synergistic task pairing. For example, several studies [11, 61, 45, 35] combine semantic segmentation with change detection from bi-temporal imagery, showing that joint learning enhances feature sharing and reduces redundancy. Another direction jointly models geometric and semantic information, such as height estimation with semantic segmentation [40, 4], demonstrating gains over single-task baselines.
Beyond specific task pairs, generalized MTL frameworks have been proposed. RSCoTr [26] performs classification, segmentation, and detection simultaneously, illustrating the potential of unified RS analysis. Large-scale datasets like SatlasPretrain [3] with multiple annotation modalities facilitate advanced MTL model development. SM3Det [33] uses a mixture-of-experts structure for multi-modal detection of horizontal and rotated bounding boxes. EarthDial [50] leverages multiple multi-task decoders to transfer knowledge across diverse tasks, enriching shared feature learning.
@@ III-B1 Data Acquisition
In this work, the dataset was curated through three sequential stages. Initially, we collected data by following the data recipes of several representative open-source vision–language models. Specifically, we adopted the description–related subsets from the instruction tuning data of VHM [46] and GeoChat [22], which cover tasks such as image classification, captioning, and visual question answering. These tasks can be further decomposed into subtasks, including modality recognition and resolution estimation. The refGeo [78] dataset was employed as the main grounding data source, while temporal multi-image data were drawn from TEOChatlas [20]. To prevent degradation of the model’s general reasoning ability during continued training, we also incorporated a subset of general-purpose data sampled from LLaVA-OneVision’s recipe [24], including chart interpretation, optical character recognition, and so on. By following these open-source data recipes, we indirectly surveyed and integrated diverse data sources.
Subsequently, we analyzed the limitations of the collected data and expanded the dataset using task-specific training set. We observed that existing RS VLMs rarely address object detection, which is crucial for fine-grained perception in RS. To fill this gap, we incorporated the DOTA-v1.0 dataset [63], thereby enriching the model’s detection-related learning capabilities.
@@ III-B2 Data Processing and Integrating
We first removed all task descriptors, such as the “[grounding]”, “[refer]”, and “[identify]” tags used in previous works [22, 46]. These descriptors tag the specific tasks. However, in open-world scenarios or novel tasks, instructions are typically expressed in natural language rather than through fixed descriptor tokens. Therefore, we replaced these descriptors with natural language prompts to better align with the real-world usage.
@@
A unified data format was further established to standardize the integration. Conversational data followed the messages structure defined by OpenAI, object detection data were formatted according to the COCO convention, and grounding data adhered to the refGeo [78] schema.
@@ III-B3 Data Loading and Weighting
We argue that the model should not rely solely on predefined prompts from the training stage. To enhance robustness, multiple agent prompts were designed for certain tasks, and one was randomly selected during training. For grounding and detection data, a unified formatting scheme was adopted. We also incorporated the JSON-based output format used in Qwen2.5-VL [2], accompanied by specific prompts, and randomly switched between standard and JSON outputs during training. In addition, a synonym replacement module was implemented to randomly substitute words with their synonyms, improving the model’s linguistic generalization. Standard data augmentation techniques, such as random resizing, were also applied to enhance multi-scale learning.
@@ III-C Dynamic Resolution Strategy
Most existing RS VLMs (such as GeoChat [22], VHM [46], and GeoGround [78]) have the only fixed square input shape (such as $336\times 336$ or $504\times 504$). For each input image, they first pad the image to a square with zeros on the right or bottom, and then resize it to the input shape. Additionally, LRS-VQA [42] and GeoLLaVA-8k [55] scale the input size to $2k\times 2k$ and $8k\times 8k$, respectively. They first cut a UHR RS image into slices of the fixed size and encode them into visual tokens. Then they prune the tokens to an amount comparable to the normal cases. In total, they pre-process the images only on a fixed shape or a small set of image shapes.
@@ III-C1 Full-scale Visual Input
The native resolution scheme in Qwen2-VL [56] inspire us to advance RS VLMs to accept inputs of arbitrary shapes. As shown in Fig. 3, let $H$ and $W$ indicate the height and width of a given RS image. $L_{patch}$ is the patch length corresponding to each visual token from the vision encoder. They first calculate the tightest shape that can wrap the input image by
@@ III-C2 Scalable Bounding Boxes
However, existing RS VLMs often suffer from a mismatch between the coordinate resolution and the input image resolution. For instance, GeoChat [22] processes images at a fixed resolution of $\times504$, but its coordinate resolution is only $\times100$, leading to a fivefold loss in localization precision and poor performance on small objects. Conversely, GeoGround [78] employs a $\times336$ input resolution but defines coordinates at a much higher $\times1000$ scale, resulting in more than half of the coordinate space being unused and excessive localization precision.